13. If Firm A operates in a perfectly competitive industry, with market price = $1,200/unit. If Firm A’s total cost function is given by TC(q)= 20q^2+ 80q + 200, find Firm A’s profit maximizing level of output.

14. Using the information from the above question: is the market in which Firm A is selling its output currently in long run equilibrium?
Answer #1

13) P = 1200$/ unit

Q = ?

Total revenue = TR = P*Q = 1200Q

Marginal revenue = d(TR)/dQ = d(1200Q)/dQ

MR = 1200$

Total cost = 2 Q​​​​​​2+80Q+200 = TC

Marginal cost = d(TC)/dQ = d(2Q2+80Q+200)/dQ

MC = 4Q+80

For profit maximization output level

MC = MR

4Q+80 = 1200

Q = 280

14) The firm is in the long run as

MR = 1200

P = 1200

MC = 4Q+80 = 4*280+80 = 1200

When the firm is having P = MR = MC than it is in long run equilibrium as they will not generate any economic profits.

Here Firm A is also in long run equilibrium as P = MC= MR